OpenStudy (anonymous):

solve by elimination. please show work, and state point of intersection. 2x+y=3 and 3x-2y=8 x+y=8 and 2x+3y=20

OpenStudy (anonymous):

x+y=8------1 2x+3y=20------2 multiply 1st equation with 2 and subtract 2nd from 1st 2x+2y= 16 -2x-3y=-20 ------------ -y=-4 y=4 now substitute y value in one of the equation x+4=8 x=8-4 x=4 so x=4,y=4

OpenStudy (anonymous):

2x+y=3 and 3x-2y=8 try it your self